OpenStudy (anonymous):

Determine algebraically whether the function is even, odd, or neither even nor odd. f(x) = 3x2 - 1

OpenStudy (freckles):

plug in -x f(-x)=? if you get f(-x)=f(x) then f is even if you get f(-x)=-f(x) then f is odd

OpenStudy (freckles):

if you get neither f(x) or -f(x) then f is neither odd or even

OpenStudy (anonymous):

like this 3(-1)^2-1

OpenStudy (anonymous):

its 2 so its even ?

OpenStudy (freckles):

yes it is even but it is even because f(-x)=3(-x)^2-1=3x^2-1=f(x) recall (-x)^2=(-1*x)^2=(-1)^2(x)^2=(1)(x)^2=x^2